OpenStudy (anonymous):

Simplify square root of 2 multiplied by the cube root of 2.

jhonyy9 (jhonyy9):

sqrt2 *\[\sqrt{2}* \sqrt[3]{2} = \]

jhonyy9 (jhonyy9):

- you can rewrite it 2^(1/2) *2^(1/3) = - hope you know that x^a *x^b = x^(a+b) so this will help you sure to solve easy this exercise